Which expression can be substituted for x in the top equation of the system to solve the system of equations by substitution? What is the solution?
-3x + 2y = 2
x + 5y = 22
Thank you, and each "/" symbol is meaning a number over a number since i can't type it that way. Thank you ! Of the 4 choices one is correct:
1)-5y + 22, (4,2)
2)-5y + 22, (2,4)
3)5y + 22, (-4,3)
4)5y + 22, (2,4)

Answer by addingup(3677)

-3x + 2y = 2
x + 5y = 22 subtract 5y from both sides: x = -5y + 22. Now use this for x.
-----------------------------------------------------------------
-3x + 2y = 2
-3(-5y + 22) + 2y = 2
15y - 66 + 2y = 2
17y - 66 = 2
17y = 68
y = 68/17
y = 4
now let's find x:
x + 5(4) = 22
x + 20 = 22
x = 2
so the answer is -5y + 22, (2,4)